OpenStudy (anonymous):

State the y-coordinate of the y-intercept for the function below. ƒ(x)=x^3-x^2-x+1

OpenStudy (saifoo.khan):

when y intercpets, x = 0 so here, ƒ(x)=x^3-x^2-x+1 ƒ(x)=(0)^3-(0)^2-0+1 ƒ(x)= +1

OpenStudy (anonymous):

so the y-coord is 1?

OpenStudy (saifoo.khan):

y intercpt is 1.

OpenStudy (saifoo.khan):

what y cordiante do u want?

OpenStudy (anonymous):

it says state the y-coord of the y-intercept. wouldnt that just be 0?

OpenStudy (saifoo.khan):

lol, thats the same thing, they are just confusing u. y intercpt = 1 y cor = (0,1)

OpenStudy (anonymous):

oh lol, thanks

OpenStudy (saifoo.khan):

hehe, welks.